OpenStudy (anonymous):

Solve for x: 4/x+4/x^2-9=3/x-3

OpenStudy (anonymous):

sorry i dont know :(

OpenStudy (anonymous):

x=9, -4

OpenStudy (anonymous):

can you elaborate???

OpenStudy (anonymous):

3/x-3 - 4/x^2-9= 4/x now, simplify the LHS and then cross multiply u will get a quadratic of form x^2-5x-36=0 find its roots, they are 9, -4 :)